. Find the slope and y-intercept of -2X + 6Y = -15
A) Slope 2 and y-intercept (0,-15) B) Slope -2 and y-intercept (0,-15) C) Slope 1/3 and y-intercept (0,-15/6) D) Slope 3 and y-intercept (0,-1/3)

Answer by checkley71(8403)

-2X+6Y=-15 or 6Y=2X-15 or Y=1/3*x-5/2 thus the slope is 1/3 and the intercept is (0,-5/2)